Can you convert this sentence into an algebraic equation and then solve it
5 more than twice a number is 29

Answer by mananth(16946)

let number be x
twice the number = 2x
5 more
= 2x+5
2x+5 =29
..
add-5
2x+5-5 = 29-5
2x=24
/2
x= 12
...
m.ananth@hotmail.ca